A 10kg box accelerates down a slope at 5m/s what is the magnitude of friction?

The incline is 15 degrees.
1
Expert's answer
2019-05-14T10:12:10-0400

A box moves with constant speed of 5 m/s. Thus,


"ma=mg \sin{15}-F_f=0"

"F_f=mg \sin{15}=(10)(9.8) \sin{15}=25\ N"